# Blending Complementary Memory Systems in Hybrid Quadratic-Linear Transformers

### v1

@@ 6 Conclusion
The quadratic transformer (KV-memory system) and linear transformer (FW-memory system) are two complementary memory systems with unique advantages. We develop a method to blend these two systems into a single hybrid memory system to leverage the strengths of both. By taking into account various types of complementarity between these two systems, we empirically evaluate our models on general language modeling, retrieval, and synthetic algorithmic tasks and reinforcement learning settings, to finally conclude that the best overall hybrid method is the one that allows KV-memory and FW-memory to operate synchronously, as opposed to an alternative that introduces a division of labor over time. This offers a novel perspective on the development of hybrid memory systems for sequence processing.
@@ A.1 Language modeling experiments
Training of 340M models using 4 H100-80GB GPUs take about 8 hours for the baseline transformer and 10 hours for DeltaNet and all the HQLT models with the window size of 64 tokens. For the 1.3 models, these numbers become 26 hours for the baseline transformer and DeltaNet, and 30 hours for all the HQLTs variants.
@@
Both the Synchronous and Delayed-Stream variants of HQLT studied in this work can directly make use of flash-attention [16] and flash-linear-attention [17] implementations (without modifying the corresponding Triton kernels) for the quadratic and linear components of the models, respectively. For Delayed-Block HQTL, we wrote a custom Triton kernel to replace intra-attention in the DeltaNet implementation of [23] by an efficient softmax attention implementation [16], and modified the backward function accordingly. Note, however, that there is still room for optimization as the intra and inter-chunk operations are currently implemented in two separate kernels (they may potentially be fused into a single kernel for further speed optimization).
@@ A.2 Synthetic algorithmic tasks
In “Modular Arithmetic (Mod Arith)”, the task is a modulo 5 addition and multiplication task (without brackets). Each symbol in an input sequence is either a number (from 00 to $4$, in the module 5 case, which is our setting), or a mathematical symbol (there are five of those: {+++, --, *, ===, eos} where the last symbol is the extra “end-of-sequence” token; which is not necessary for the modulo 5 case but included by convention). Here not only the output, but also the numbers in the sequences, are drawn between 00 to $4$. This makes the total vocabulary size of $10$ with a chance level accuracy of 20%.
@@
We train with a batch size of 1024 for 20,000 steps. We search for the best learning rate among {$5e^{-3}$, $1e^{-3}$, $5e^{-4}$, $1e^{-4}$} (the only difference compared to Grazzi et al. [8] is that this list includes $5e^{-3}$ instead of $1e^{-2}$), each with three seeds. We directly measure the validation accuracy to determine the best configuration. In the main result table, Table 3, we report the best/max result among the seeds for the best configuration, while Table 6 shows variability among seeds. This is a standard practice in formal language recognition tasks [65, 9]. We train with sequence lengths from 3 to 40, and validate on sequences of lengths from 40 to 256.
Each training run on a single H100 takes about 70 min.
@@ Appendix B Further discussion
B.1 Practical computational costs of increasing the window size
In Sec. 4.3/Table 4, we demonstrated how increasing the quadratic attention window size in HQLTs improves their retrieval abilities. Here we discuss the computational costs for such improvements. In practice, within the range of window sizes discussed here, the actual computational cost is negligible, since during inference with a batch size of 1, short-window attention can be efficiently computed with a few/two steps of parallelizable matrix multiplication on a GPU, regardless of whether the window size is 64 or 1024. Training time is also only minimally affected by the increased window size as training is parallelized over the sequence elements. However, the total state size (which has a direct impact on the memory requirement) increases as a function of the window size $S$ as $S(d_{\text{in}}+d_{\text{out}})+d_{\text{out}}*d_{\text{in}}$ per layer and per head.
B.2 Positional Encodings in KV-memory
All the HQLT models presented in the main text use the RoPE positional encodings [61] in KV-memory. However, it is also known that multi-layer self-attention can process sequences without explicit positional encodings [67, 68, 69], and such a “No Pos” approach has certain generalization benefits [65, 70]. Here we provide an ablation study on the positional encoding in KV-memory. We compare the Synchronous HQLT models with and without RoPE for the 340M models with a window size of 64 or 1024, as well as for the 1.3B model with a window size of 64.

### v2

@@ 6 Conclusion
The quadratic transformer (KV-memory system) and linear transformer (FW-memory system) are two complementary memory systems with unique advantages. We develop a method to blend these two systems into a single hybrid memory system to leverage the strengths of both. By taking into account various types of complementarity between these two systems—complexity, context length, retrieval precision and expressivity, we empirically evaluate our models on general language modeling, retrieval, and synthetic algorithmic tasks and reinforcement learning settings, to finally conclude that the best overall hybrid method is the one that allows KV-memory and FW-memory to operate synchronously, as opposed to alternatives that introduce a division of labor over time. This offers a novel perspective on the development of hybrid memory systems for sequence processing.
@@ A.1 Language modeling experiments
Training time. Training of 340M models using 4 H100-80GB GPUs take about 8 hours for the baseline transformer and 10 hours for DeltaNet and all the HQLT models with the window size from 64 to 1024 tokens. For the 1.3B models, these numbers become 26 hours for the baseline transformer and DeltaNet, and 30 hours for all the HQLTs variants. Given that the training for the 1.3B models uses 16B tokens, training throughputs are 170K tokens/second and 148K tokens/second for the baseline Transformer/DeltaNet and HQLT, respectively.
@@
Both the Synchronous and Delayed-Stream variants of HQLT studied in this work can directly make use of flash-attention [23] and flash-linear-attention [24] implementations (without modifying the corresponding Triton kernels) for the quadratic and linear components of the models, respectively. For Delayed-Block HQLT, we wrote a custom Triton kernel to replace intra-attention in the DeltaNet implementation of [29] by an efficient softmax attention implementation [23], and modified the backward function accordingly. Note, however, that there is still room for optimization as the intra and inter-chunk operations are currently implemented in two separate kernels (they may potentially be fused into a single kernel for further speed optimization).
@@ A.2 Synthetic algorithmic tasks
In “Modular Arithmetic (Mod Arith)”, the task is a modulo 5 addition and multiplication task (without brackets). Each symbol in an input sequence is either a number (from 0 to $4$, in the module 5 case, which is our setting), or a mathematical symbol (there are five of those: {++, -, *, ==, eos} where the last symbol is the extra “end-of-sequence” token; which is not necessary for the modulo 5 case but included by convention). Here not only the output, but also the numbers in the sequences, are drawn between 0 to $4$. This makes the total vocabulary size of $10$ with a chance level accuracy of 20%.
@@
We train with a batch size of 1024 for 20,000 steps. We search for the best learning rate among $5e^{-3}$, $1e^{-3}$, $5e^{-4}$, $1e^{-4}$} (the only difference compared to Grazzi et al. [7] is that this list includes $5e^{-3}$ instead of $1e^{-2}$), each with three seeds. We directly measure the validation accuracy to determine the best configuration. In the main result table, Table 3, we report the best/max result among the seeds for the best configuration, while Table 6 shows variability among seeds. This is a standard practice in formal language recognition tasks [12, 10]. We train with sequence lengths from 3 to 40, and validate on sequences of lengths from 40 to 256. Each training run on a single H100 takes about 70 min.
@@ Appendix B Further discussion
B.1 Further clarifications for the Delayed-Chunk variant
Delayed-Streaming HQLT and Delayed-Chunk HQLT are not equivalent (i.e., their outputs are different for the same input) because the classic equivalence between recurrent vs. chunk mode computation in linear attention does not hold when softmax is applied within the intra-chunk attention. To be more specific, the Delayed-Streaming variant has a sliding window attention with a stride of 1, i.e., at every time step, the oldest token from the KV-memory window is fed to the FW-memory, and a new token enters the KV-memory window; as a consequence, the memory content of FW-memory is updated at every time step, whereas in the Delayed-Chunk variant, the sequence is processed chunk-by-chunk, i.e., the content of FW-memory is only updated at the chunk-boundaries and remains constant while the system is processing the current chunk. Within a chunk, the content of KV-memory is updated at every time step by adding the new element in the chunk; its content is reset (to empty) at the chunk boundaries. Figure 3 provides an illustration. Delayed-Chunk HQLT is naturally compatible with efficient training: we can apply flash-linear-attention for inter-chunk attention computation, and flash-attention to the intra-chunk attention with softmax.
Intuitively, Delayed-Streaming HQLT may be argued to be a better approach as the KV-memory always makes use of the full window size (except at the very beginning of the sequence), but we also included Delayed-Chunk which is not only a natural extension of chunk-wise parallel form of linear attention, but was used in prior work by Munkhdalai et al. [22].
Figure 3: An illustration for the Chunk-Delayed variant of Hybrid Quadratic-Linear Transformers (HQLTs). Here the chunk size is 4. KV-memory only has access to the tokens within the current chunk (i.e., the last three tokens here), whereas FW-memory does not contain any information about the current chunk.
B.2 Practical computational costs of increasing the window size
In Sec. 4.3/Table 4, we demonstrated how increasing the quadratic attention window size in HQLTs improves their retrieval abilities. Here we discuss the computational costs for such improvements. In practice, within the range of window sizes discussed here, the extra computational cost is negligible, since during inference with a batch size of 1, short-window attention is typically memory-bound rather than compute-bound and can be efficiently computed in a few parallelizable matrix multiplication steps on a GPU, regardless of whether the window size is 64 or 1024. Training time is also only minimally affected by the increased window size as training is parallelized over the sequence elements. However, the total state size (which has a direct impact on the memory requirement) increases as a function of the window size $S$ as $S(d_{\text{in}}+d_{\text{out}})+d_{\text{out}}*d_{\text{in}}$ per layer and per head.
@@ B.2 Practical computational costs of increasing the window size
B.3 Positional Encodings in KV-memory
All the HQLT models presented in the main text use the RoPE positional encodings [68] in KV-memory. However, it is also known that multi-layer self-attention can process sequences without explicit positional encodings [73, 74, 75], and such a “No Pos” approach has certain generalization benefits [12, 76]. Here we provide an ablation study on the positional encoding in KV-memory.
We compare the Synchronous HQLT models with and without RoPE for the 340M models with a window size of 64 or 1024, as well as for the 1.3B model with a window size of 64.